# Neural Fields with Thermal Activations for Arbitrary-Scale Super-Resolution

### v2

@@ 4.3 Ablation Studies
In Tab. 4, we ablate individual components and design choices of our method to understand their contributions to overall performance. These comparisons use Thera-M with EDSR-baseline backbone, but are also representative of other settings.
Single scale training. We ran three experiments using a single scale ($\times$2, $\times$3, or $\times$4) to test how this affects scale generalization. $\kappa$ was fixed for these experiments as multi-scale training is required to optimize it. As expected, we observe equal or even superior performance compared of single-scale training when tested at the training scale (marked in yellow in Tab. 4), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
Trainable $\bm{\kappa}$. Fixing $\kappa$ at the theoretically derived value (Eq. 5) leads to a small drop in performance. This suggests that there remain effects that are not accounted for by our proposed observation model, but that these are very minor.
Geometric self-ensemble. In line with previous work [42, 24, 43] we see a notable performance boost with geometric self-ensembling (GSE). Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
Total variation. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
No thermal activations. We replace thermal activations (Eq. 4) with standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
Shared components. We observe a noticeable performance drop, at all scales, if we make the components $\mathbf{W}_{1}$ an output of the hypernetwork $\Psi$. Despite the greater flexibility afforded by individual per-field components, it seems challenging to learn them jointly with other, correlated parameters.
Table 4: Ablation experiments with Thera-M using the EDSR-baseline backbone.
Method In-distribution Out-of-distribution
@@
Thera-M 34.86 31.21 29.22 26.94 22.33 20.60
$\times$2 only, fixed $\kappa$ 34.98 30.96 28.81 26.40 21.93 20.30
$\times$3 only, fixed $\kappa$ 34.71 31.22 29.22 26.88 22.24 20.53
$\times$4 only, fixed $\kappa$ 34.44 31.14 29.22 26.95 22.30 20.58
Fixed $\kappa$ 34.85 31.20 29.22 26.93 22.30 20.58
No GSE 34.79 31.14 29.16 26.88 22.28 20.56
No TV prior 34.87 31.21 29.21 26.86 20.82 19.12
ReLU instead of $\xi$ 34.80 31.12 29.13 26.87 22.28 20.58
Non-shared comps. 34.80 31.13 29.14 26.87 22.28 20.57
@@ 4.4 Limitations and Future Work
Thermal activations as introduced in this paper, and by extension our proposed ASSR method, come with relatively strict architectural requirements that only allow for a single hidden layer in the neural field. While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder. We have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, and we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work. This could result in even higher parameter efficiency, and potentially better generalization. It may also be possible to construct similar formulations that do not require the usage of a hypernetwork, which could enable more streamlined training.
Relative to other methods, Thera’s performance slightly drops for $\times$2 super-resolution when compared to higher scaling factors. We speculate that at small scaling factors Thera’s isotropic, continuous field representation is increasingly challenged by anisotropic effects from the oriented, discrete pixel grid. Further explorations are necessary to better understand this transition regime and perhaps come up with hybrid representations (which could potentially benefit also other methods at the interface of gridded and implicit representations like Instant-NGP [33]).
Finally, we expect that more advanced priors than TV could be even more effective at regularizing $\Phi$. Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR performance for all scaling factors.
@@ 5 Conclusion
We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations. Our method, Thera, implicitly describes an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) analytically, with negligible overhead. Our experimental evaluation shows that Thera outperforms competing methods in most cases, while at the same time being more parameter efficient, and offering theoretical guarantees. We are convinced that Thera-style representations could also benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful observation models like our Gaussian PSFs.
Appendix 0.A Theory
0.A.1 Preliminaries
As was described in Sec. 3.1, the idea underlying our neural field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
@@ 0.A.1 Preliminaries
0.A.2 Thermal Diffusivity Coefficient
@@
For the Thera formulation presented in Sec. 3.2, we want $\Phi$ to contain a single pixel at $t=1$. This is the pixel from the low-resolution input which will become $SR^{2}$ pixels for super-resolution with a scaling factor of $SR$. Therefore we initialize $\kappa$ with
@@ 0.A.3 Relationship Between $t$ and $s$
For instance, if we want to upsample the image by a factor of 2, we should use
$t=0.5^{2}=0.25.$ (17)
Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling. This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
0.A.4 Other Filters
In theory, the formulation presented in Sec. 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely. Gaussian filters are an obvious choice since they are often used as anti-aliasing filters and since they are fully defined by a single parameter, the standard deviation. Initial explorations of a sharp low-pass filter that completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely due to the associated effect on gradients during training. It remains an open question whether more complex filters (e.g., Butterworth filters) would improve the current formulation in any way. For quantitative evaluations, this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios, this may be desirable.
0.A.5 Initialization of Components
@@
Appendix 0.B Additional Qualitative Results
Figures 9 and 10 contain additional qualitative results for lower and higher scaling factors, respectively. Again, we see that Thera can reconstruct the targets more faithfully. This is particularly (but not exclusively) visible in repetitive structural elements. It seems evident that Thera’s inbuilt Gaussian PSF and the sinusoidal basis functions contribute to this. Furthermore, we provide a showcase of the continuous upsampling capabilities of our method in Fig. 6.
Figure 6: Showcase of multiscale upsampling using Thera-L with a SwinIR [23] backbone, shown with non-integer scaling factors.
Appendix 0.C Additional Quantitative Results
Table 5 contains SSIM metrics for Thera on DIV2K, which complement the PSNR values reported in Tab. 1. We observe strong performance by all Thera variants, although overall there is relatively little variance across all methods using this metric. Furthermore, Tab. 6 contains quantitative evaluations of Thera which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$). Finally, in Fig. 7 we compare the number of additional parameters and PSNR values for various methods and upsampling factors on DIV2K.
Figure 7: Plots depicting the number of additional parameters vs. PSNR on DIV2K using the EDSR-baseline backbone for the ×$\times 2$ (left), ×$\times 3$ (center) and ×$\times 4$ (right) upsampling tasks.
Table 5: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the SwinIR [23] backbone for all models.
In-distribution Out-of-distribution
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$$\times 12$$\times 18$$\times 24$$\times 30$
LIIF [11] .942 .883 .830 .753 .702 .641 .594 .568 .551
LTE [22] .942 .883 .832 .755 .704 .643 .595 .569 .552
CiaoSR [8] .942 .883 .832 .756 .706 .645 .598 .572 .551
Thera-S .943 .884 .833 .757 .706 .644 .596 .569 .552
Thera-M .943 .885 .834 .758 .707 .645 .596 .569 .552
Thera-L .942 .885 .834 .759 .708 .646 .597 .570 .552
Table 6: Evaluation on common benchmark datasets for out-of-distribution scale factors, using the SwinIR backbone. The numbers represent PSNR in dB, calculated on the luminance (Y) channel of the YCbCr representation following previous work.
@@ Appendix 0.C Additional Quantitative Results
$\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12 $\times$6 $\times$8 $\times$12
MetaSR 29.09 27.02 24.82 26.58 25.09 23.33 25.94 24.87 23.59 24.16 22.75 21.31 27.29 24.96 22.35
LIIF 29.46 27.36 26.82 25.34 26.07 25.01 24.59 23.14 27.69 25.28
LTE 29.50 27.35 26.86 25.42 26.09 25.03 24.62 23.17 27.83 25.42
CUF 29.57 26.91 26.11 24.69
CiaoSR 29.62 27.45 24.96 26.88 25.42 23.38 26.13 25.07 23.68 24.84 23.34 21.60 28.01 25.61 22.79
CLIT 29.69 27.62 26.94 25.55 26.15 25.09 24.77 23.33
Thera-S (ours) 29.45 27.33 25.00 26.85 25.40 23.40 26.11 25.05 23.69 24.68 23.23 21.55 27.94 25.49 22.73
Thera-M (ours) 29.47 27.43 25.07 26.90 25.41 23.40 26.13 25.07 23.69 24.70 23.25 21.57 28.05 25.59 22.79
Thera-L (ours) 29.48 27.46 25.09 26.94 25.45 23.42 26.14 25.08 23.71 24.79 23.32 21.62 28.13 25.69 22.85
Appendix 0.D Hypernetwork Architecture
In Fig. 8, we show in more detail the hypernetworks used for the variants Thera-S, -M, and -L. The backbone is the same across the three variants, and extracts 64-dimensional feature maps from the input image. It is then followed by a number of (residual) ConvNeXt [26] blocks. Each block consists of (i) a $3\times 3$ convolution that maps to $d$-dimensional features, (ii) a layer normalization [2] operation, (iii) a $1\times 1$ convolution with $4d$ filters, (iv) a GELU [15] non-linearity, and (v) another $1\times 1$ convolution with $d$ filters. Note that we decreased the kernel size of each block’s first convolution from $7\times 7$ – as originally used by Liu et al. [26] – to $3\times 3$ to reduce the number of parameters. Furthermore, note that in Thera-L, the first convolution per block uses $d$ groups (effectively implementing a channel-wise convolution), while for the two smaller variants we used full convolutions, i.e., a single group. In our implementation, Thera-S uses zero ConvNeXt blocks, Thera-M uses 5 blocks with $d=64$, and Thera-L uses 6 blocks with $d=64$ followed by 7 blocks with $d=96$ and 3 blocks with $d=128$. For Thera-L, we add further projection blocks between ConvNeXt blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation.
Figure 8: Schematics of the hypernetworks for a) Thera-S, b) Thera-M and c) Thera-L.
@@ Appendix 0.D Hypernetwork Architecture
Finally, a $1\times 1$ convolution maps pixel-wise features into field parameters, as done in SIREN [40]. There are 128 field parameters for Thera-S and 2048 for the larger variants Thera-M and Thera-L. Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork. The total parameter count of the hypernetwork is 8,192, 0.48 M, and 1.41 M for the three variants, respectively. To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters. We highlight that these do not come from the hypernetwork.
Appendix 0.E Replication of Other Methods

### v3

@@ 4.2 Ablation Studies
Experiment In-distribution Out-of-distribution
@@
Thera Plus 34.89 31.22 29.24 26.96 22.34 20.61
$\times$2 only, fixed $\kappa$ 35.01 31.00 28.84 26.47 21.97 20.33
$\times$3 only, fixed $\kappa$ 34.74 31.25 29.25 26.91 22.25 20.54
$\times$4 only, fixed $\kappa$ 34.47 31.17 29.24 26.97 22.32 20.59
Fixed $\kappa$ 34.86 31.22 29.24 26.95 22.31 20.58
No GSE 34.81 31.15 29.18 26.91 22.29 20.57
No TV prior 34.89 31.23 29.24 26.87 20.42 18.68
ReLU instead $\xi$ 34.80 31.11 29.13 26.87 22.28 20.57
Predicted comps. 34.90 31.23 29.25 26.97 22.34 20.61
Table 3: Ablation study using Thera Plus (w/ EDSR-baseline)
In Table 3, we ablate individual components and design choices of our method to understand their contributions to overall performance. The comparisons use Thera Plus with EDSR backbone, and are representative of all variants.
Single scale training. We run three experiments using a single scale ($\times$2, $\times$3, $\times$4) to test how this affects scale generalization. $\kappa$ was fixed at the theoretically derived value for these experiments, as multi-scale training is required to optimize it. As expected, we observe equal or even superior performance of single-scale training when tested at the training scale (marked in yellow in Tab. 3), but a significant drop compared to the default multi-scale version when generalizing to other scaling factors.
Trainable $\bm{\kappa}$. Fixing $\kappa$ at the theoretically derived value (Eq. 3) leads to a small drop in performance. This suggests that there remain effects that are not accounted for by our proposed observation model, albeit very minor.
Geometric self-ensemble. In line with previous work [43, 26, 44] we see a notable performance boost with geometric self-ensembling. Note, though, if an application prioritizes inference speed over quality this add-on can be disabled at test-time without re-training the network.
Total variation prior. The regularizer has a negligible effect for in-domain scaling factors, but performance degrades significantly without it for out-of-distribution scales.
Thermal activations. We replace thermal activations (Eq. 2) with standard ReLU activations. What remains is only the hypernetwork controlling the parameters of the local fields. A consistent loss in performance shows the impact of the proposed thermal activation underlying our multi-scale representation.
Shared components. Predicting $\bm{W}_{1}$ along with $\bm{b}_{1}$ and $\bm{W}_{2}$ leads to negligible gains. This comes at the cost of doubling the amount of field parameters that the hypernetwork predicts. Thus, Thera uses a shared, global frequency bank.
@@ 4.3 Limitations and Future Work
Neural heat fields as introduced in this paper, and by extension Thera, come with relatively strict architectural requirements that currently only allow for a single hidden layer in the neural field. While this can be beneficial from a computational standpoint, it limits hierarchical feature learning and potentially makes modeling of complex non-linear relations harder than necessary. Nonetheless, as our experiments show, the current neural heat field architecture does easily have enough capacity to model local, subpixel information for the scaling factor range discussed in this paper. We have compensated for the relatively less expressive fields with a higher-capacity hypernetwork, and we speculate that there may be ways to extend the signal-theoretic guarantees of Thera to multi-layer architectures in future work. This could result in even higher parameter efficiency, and potentially better generalization. We also expect that more advanced priors than TV could be even more effective at regularizing $\Phi$. Priors at $t=0$, made possible by Thera, have the potential to regularize the continuous signal itself, and therefore improve SR quality for all scaling factors.
@@ 5 Conclusion
We have developed a novel paradigm for arbitrary-scale super-resolution by combining traditional signals theory with modern implicit neural representations. Our proposed neural heat fields implicitly describe an image as a combination of sinusoidal components, which can be selectively modulated according to their frequency to perform Gaussian filtering (anti-aliasing) between scales analytically, with negligible overhead. Our experimental evaluation shows that Thera, our ASR method based on neural heat fields, consistently outperforms competing methods. At the same time, it is more parameter-efficient and offers theoretical guarantees w.r.t. aliasing. We believe that Thera-style representations could benefit other computer vision tasks and hope to inspire further research into neural methods that integrate physically meaningful and theoretically grounded observation models.
Appendix A Theory
A.1 Preliminaries
As was described in Section 3.1, the idea underlying our neural heat field with thermal activations is to formulate a neural field $\Phi(\mathbf{x},t)$, with $\mathbf{x}$ being the 2-dimensional spatial coordinates $(x_{1},x_{2})$, such that $\Phi$ follows the heat equation:
@@ A.1 Preliminaries
A.2 Thermal Diffusivity Coefficient
@@
For our proposed Thera formulation, we want $\Phi$ to contain a single pixel at $t=1$. This is the pixel from the low-resolution input which will become $SR^{2}$ pixels for super-resolution with a scaling factor of $SR$. Therefore we initialize $\kappa$ with
@@ A.3 Relationship Between t and s
For instance, if we want to upsample the image by a factor of 2, we should use $t=0.5^{2}=0.25$. Thus, $10<t<$ 1 refers to super-resolution, while $t>1$ refers to downsampling. This is intuitive: As $t$ grows, the image becomes blurrier (the Gaussian kernel gets wider), which corresponds to stronger low-pass filters and therefore lower sampling rates.
Image domain
Refer to caption Refer to caption Refer to caption Refer to caption
Original $\times$1/4 AA $\times$1/8 AA $\times$1/16 AA
Neural field
Refer to caption Refer to caption Refer to caption Refer to caption
$\Phi(\mathbf{x},1)$ $\Phi(\mathbf{x},4^{2})$ $\Phi(\mathbf{x},8^{2})$ $\Phi(\mathbf{x},16^{2})$
Figure 6: Comparison between a theoretical anti-aliasing filter (top row) and anti-aliasing with neural heat fields (bottom row), which is computed without convolutions or over-sampling. The heat field was supervised at $\Phi(\mathbf{x},1)$ only.
Figure 6 shows an example where we fit a neural heat field at $t=1$ to the image. After training, any low-pass filtered version of the image can be generated by setting $t$ according to Eq. 4. We emphasize that: (i) Computing these filtered images requires no over-sampling or convolutions; (ii) The computational cost does not depend on the size of the blur kernel or on $t$; (iii) Given $\Phi(\mathbf{x},t_{0})$, the filtered versions $\Phi(\mathbf{x},t)$ are known for any $t\geq t_{0}$.
$t=0$
Refer to caption Refer to caption Refer to caption Refer to caption
$t=1$
Refer to caption Refer to caption Refer to caption Refer to caption
a) b) c) d)
Figure 7: Example situations where aliasing would occur without the suppression of high frequencies as modeled by neural heat fields. Sampling the center location at $t=0$ would not be representative of the pixel’s footprint.
Figure 8: Blurring an image before a point-wise sample is equivalent to observing with PSF equivalent to the blur kernel.
In Fig. 7, we show four local neural heat fields in which aliasing would occur without the anti-aliasing mechanism modeled by thermal activations. Note that aliasing is not always as obvious to the eye as Moiré patterns: in the cases shown in the figure, it would simply mean that sampling the center location at $t=0$ would not be representative of the pixel’s footprint. Figure 8 further illustrates how such blur is equivalent to a scale-appropriate anti-aliasing filter.
A.4 Other Filters
In theory, the formulation presented in Section 3.1 allows us to use any low-pass filter we want, since we can modulate different components freely. Gaussian filters are an obvious choice since they are often used as anti-aliasing filters, and since they are fully defined by a single parameter, the standard deviation. Initial explorations of a sharp low-pass filter that completely removes components above $f_{\text{Nyquist}}$ led to a performance reduction, likely due to the associated effect on gradients during training. It remains an open question whether more complex filters (e.g., Butterworth) would improve the current formulation in any way. For quantitative evaluations, this is unlikely, since the downsampling operations use Gaussian anti-aliasing, but in real-world applications or other scenarios, this may be desirable.
A.5 Initialization of Components
@@
Appendix B Continuous Upsampling Example
In Fig. 9, we provide a practical showcase of the continuous upsampling capabilities of our method.
Figure 9: Showcase of multiscale upsampling using Thera Pro with a RDN [54] backbone, shown with non-integer scaling factors.
Appendix C Additional Quantitative Results
C.1 Further Metrics
In-distribution OOD
×$\times 2$$\times 3$$\times 4$$\times 6$$\times 8$
LIIF [10] .934 .866 .804 .704 .636
LTE [24] .936 .868 .807 .707 .639
CiaoSR [7] .938 .871 .814 .718 .651
SRNO [45] .937 .873 .819 .738 .685
MSIT [56] .942 .882 .829 .751 .700
Thera Pro .943 .884 .833 .757 .706
Table 4: SSIM scores (higher is better) for several methods and scaling factors evaluated on the (hold-out) DIV2K validation set. We use the RDN backbone for all models. Some methods did not provide checkpoints at the time of writing, see Appendix F.
@@ C.1 Further Metrics
$\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8 $\times$6 $\times$8
MetaSR 29.04 29.96 26.51 24.97 25.90 24.83 23.99 22.59
LIIF 29.15 27.14 26.64 25.15 25.98 24.91 24.20 22.79 27.33 25.04
LTE 29.32 27.26 26.71 25.16 26.01 24.95 24.28 22.88 27.49 25.12
CUF 29.27 26.74 26.03 24.32
CiaoSR 29.46 27.36 26.79 25.28 26.07 25.00 24.58 23.13 27.70 25.40
CLIT 29.39 27.34 26.83 25.35 26.07 25.00 24.43 23.03
SRNO 29.38 27.28 26.76 25.26 26.04 24.99 24.43 23.02 27.66 25.31
MSIT 29.34 27.29 26.75 25.26 26.05 24.98 24.43 22.99 27.61 25.26
Thera Air 29.31 27.25 26.76 25.27 26.05 24.99 24.39 23.00 27.70 25.30
Thera Plus 29.31 27.29 26.80 25.32 26.07 25.00 24.45 23.04 27.80 25.41
Thera Pro 29.51 27.34 26.90 25.38 26.12 25.04 24.70 23.24 27.94 25.49
Table 5: PSNR (Y channel) on common benchmark datasets for out-of-distribution scale factors, with an RDN [54] backbone. For some methods, code and/or checkpoints were not publicly avalibale, see Appendix F.
Table 4 shows SSIM metrics on the DIV2K validation set, which complement the PSNR values reported in Table 1. We use the SSIM implementation from torchmetrics [13]. We observe strong performance of Thera, although overall there is relatively little variance across all methods using this metric.
Furthermore, in Table 5 we show quantitative evaluations which are out of distribution both in terms of data (benchmark datasets) and in terms of scaling factors (above ×$\times 4$).
C.2 Parameter Efficiency
In Fig. 10 we compare the number of additional parameters and PSNR values for various methods and individual upsampling factors on the DIV2K validation set.
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$34.734.734.734.734.834.834.834.834.934.934.934.93535353535.135.135.135.135.235.235.235.2Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×2 Scaling Factor
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$3131313131.131.131.131.131.231.231.231.231.331.331.331.331.431.431.431.431.531.531.531.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×3 Scaling Factor
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$2929292929.129.129.129.129.229.229.229.229.329.329.329.329.429.429.429.429.529.529.529.5Thera AirThera PlusThera Pro(Ours)LIIFLTECUFCiaoSRCLITSRNOMSITModel Size (M parameters)PSNR (dB)×4 Scaling Factor
Figure 10: Comparison of ASR methods for different scaling factors (×2, ×3, and ×4) on DIV2K. Thera consistently achieves better performance at lower parameter counts across all scaling factors.
Appendix D Hypernetwork Architecture
In our implementation, Thera Air uses no feature refinement blocks on top of the backbone except a $1\times 1$ convolution mapping pixel-wise features into field parameters, as done in SIREN [41]. Thera Plus uses 6 ConvNeXt [29] blocks with $d=64$ followed by 7 ConvNeXt blocks with $d=96$ and 3 ConvNeXt blocks with $d=128$, prior to the final mapping layer. Projection blocks are added between blocks with different $d$, which consist of a layer normalization and a $1\times 1$ convolution operation. For Thera Pro, two SwinIR [25] blocks are used with depth 7 and 6, respectively, and 6 attention heads per layer.
@@ Appendix D Hypernetwork Architecture
There are 128 field parameters for Thera Air and 2048 for the larger variants Thera Plus and Thera Pro. These numbers are computed as follows: Let $c$ be the number of components used (32, 512, and 512 respectively), then we need $c$ field parameters indicating phase shift and $3c$ field parameters for the linear mapping between components and RGB channels, resulting in a total of $4c$ field parameters produced by the hypernetwork. The total parameter count of the hypernetwork is 8,192, 1.41 M, and 4.63 M for the three variants, respectively. To this we have to add the parameters of the components themselves (i.e., the mapping from coordinate space to thermal activation arguments defined by $\mathbf{W}_{1}$) that are defined in a global, learnable frequency bank, adding further $2c$ parameters. We highlight that these do not come from the hypernetwork.
Appendix E Real-World Optical Zoom Data
To evaluate the effectiveness of our approach on real-world continuous optical zoom data, we conducted experiments using the COZ dataset [14]. We introduce Thera++, which combines two components: (1) Thera Plus (EDSR) trained on the COZ dataset, and (2) a lightweight spatial transformer network (STN) [21] that estimates just 6 parameters per image to correct domain-specific affine distortions. Despite its name, the STN is not a transformer network in the modern sense that employs attention mechanisms; rather, it is an image model block that explicitly allows the spatial manipulation of data within a convolutional neural network.
The STN component follows a standard architecture, consisting of a simple convolutional localization network with adaptive pooling to handle variable input sizes. The network processes the input image along with the scale factor and outputs six parameters of a 2D affine transformation ($xy$-translation, anisotropic $xy$-scale, rotation and shear). With approximately 10K parameters, this lightweight component efficiently corrects for geometric distortions while adding minimal computational overhead.
Thera++ addresses an important limitation of the COZ dataset. When looking at the data it becomes obvious that COZ samples are not perfectly aligned, resulting in $xy$-jitter between images. Additionally, dynamic objects like people, dust, leaves, and moving shadows appear inconsistently across images of the same scene, significantly increasing the noise level. These challenges make super-resolution particularly difficult for this dataset. However, Thera++ outperforms previous state-of-the-art methods, including LMI, across all scaling factors, highlighting its applicability under real-world imaging conditions, see Tab. 6.
Method In-distribution Out-of-distribution
$\times$2 $\times$3 $\times$4 $\times$5 $\times$6
MetaSR [18] 28.70 26.55 25.17 24.31 23.25
LIIF [10] 28.72 26.61 25.16 24.32 23.23
LTE [24] 28.67 26.55 25.15 24.37 23.26
SRNO [45] 28.73 26.59 25.15 24.31 23.25
LIT [9] 28.74 26.58 25.15 24.35 23.19
LMI [14] 28.86 26.66 25.22 24.39 23.29
Thera++ 29.06 26.84 25.45 24.49 23.46
Table 6: Results (PSNR in dB) on the COZ test set.
Appendix F Reproducibility of Existing Methods